### v1

@@ 1 Introduction
Artificial intelligence (AI) is increasingly transforming manufacturing, where computer vision technologies play central roles in many tasks such as monitoring processes, detecting defects, and ensuring quality [mozaffar2025tunes, yazdipaz2025robust, safavigerdini2025automated, dashti2025functional]. However, most vision-based systems today depend on large labeled datasets that are built for specific tasks and environments. This reliance makes them hard to adapt in dynamic manufacturing settings, since collecting and labeling data is both time-consuming and labor-intensive. Many industrial environments are low light or enclosed, which reduces the reliability of standard optical cameras. Thermal infrared (IR) imaging offers a strong alternative because it does not depend on ambient light and can capture useful contrast from temperature differences between machine surfaces, tools, and parts. Figure 1 illustrates this advantage.
@@
Zero-shot learning (ZSL) offers a promising alternative that enables models to generalize to unseen tasks without the need for retraining [xian2017zero, aghdam2025actalign]. ZSL methods have also been explored in manufacturing. Hua et al. [hua2022zero] introduced a framework for tool wear prediction under non-stationary machining environments. They combined deep learning with causal inference to extract invariant features across conditions. A tensor-based method for zero-shot fault diagnosis in smart process manufacturing was also developed to evolve dual prototypes to align sample and attribute representations [ren2024zero]. They improved generalization to unseen faults across multiple datasets. Huang et al. [huang2023simple] presented a simple framework for generalized zero-shot fault diagnosis, using binary classifiers with semantic correction to reduce bias in fault attributes. These studies highlight the strength of ZSL in manufacturing because collecting and labeling large datasets is both costly and often impractical.
@@
As discussed above, VLMs have recently shown promise in manufacturing, but their use in this domain is still at an early stage [zhao2025industrial]. A key limitation for utilizing these models in manufacturing tasks is that they are trained almost entirely on natural RGB images [faysal2024nmformer]. This makes it unclear how well they can transfer to other modalities such as thermal infrared imagery [borazjani2025multi]. This is an important gap because thermal cameras are already common in low-light or enclosed industrial environments where RGB cameras are unreliable. Some studies in other domains have adapted vision foundation models to infrared data, but no prior work has systematically investigated zero-shot classification with VLMs on thermal IR data in manufacturing tasks.
@@
This work introduces VLM-IRIS (Vision–Language Models for InfraRed Industrial Sensing), a zero-shot framework for adapting VLMs to thermal infrared imagery. It combines the strengths of zero-shot VLMs and infrared imaging to enable reliable object presence detection without any model retraining. The framework converts the infrared images into RGB representations suitable for VLMs to bridge the modality gap. It also utilizes a centroid-based prompting strategy, which multiple descriptive sentences are averaged to create stable text embeddings that reduce sensitivity to prompt wording in order to capture the highest similarity between image-text pairs in the VLMs’ semantic shared space. The best-performing configuration of the framework combines CLIP with magma preprocessing and centroid-based prompting to achieve strong accuracy in zero-shot classification. To further validate the robustness of VLM-IRIS, we also evaluate alternative settings across different preprocessing pipelines, prompting strategies, and dataset conditions.
@@ 2.3 Embeddings and Classification via Cosine Similarity
Рис. 3: Example of image-text pairs in shared embedding space. Infrared images of an empty and an occupied 3D printer are projected alongside their textual prompts, showing how the alignment of vectors reflects semantic similarity.
After preprocessing the infrared images and defining the prompt banks, a pretrained VLM was utilized to project both the RGB images and the prompts into a shared multimodal embedding space. The model employs two separate encoders: one for images and one for text. The image encoder transforms each infrared image into a feature vector while the text encoder maps each prompt into the same space. All vectors were normalized to unit length for evaluation based on their directions. Figure 3 illustrates this process in a simplified two-dimensional embedding space.
@@ 3 Experimental Setup
Рис. 4: Experimental setup featuring a FLIR Boson thermal camera and a Prusa MK3S 3D printer on the table.
@@ 3.1 Data Collection
The Prusa MK3S 3D printer was used to test the framework as a representative manufacturing platform. A FLIR Boson thermal infrared camera was mounted above the printer’s build plate to capture top-down images of the heated bed after printing. Figure 4 shows the experimental setup that was used for data collection. All test parts were fabricated using PETG filament with the nozzle temperature set to $\sim$230 °C and the bed temperature maintained between 85–110 °C. The collected dataset consists of two classes: object present and object absent. To evaluate the effect of thermal contrast between the surface and the printed part on predictions, data points were captured under a hot build plate ($\sim$85°C) and a room temperature build plate ($\sim$34°C). In total, 200 images were captured, with a balanced distribution across the two categories and temperature conditions. To further evaluate generalization, objects with different geometries, shapes, and sizes were printed. Figure 5 shows examples of the printed objects used in this study.
Рис. 5: A few 3D printed parts during the experiments. This figure presents the variety of the shapes and geometries of the printed parts.
@@ 3.2 Preprocessing Variants
In grayscale mapping, the single thermal intensity channel is normalized and replicated across three channels to preserve the true temperature contrast without adding artificial colors. This method serves as a baseline to test whether VLMs can interpret thermal data that is visually simple yet physically accurate. Magma and viridis colormaps were chosen because they are widely used in scientific visualizations. These colormaps transform intensity values into structured RGB color distributions, where pixel intensity values are mapped to colors. Such mappings can introduce artificial edges and gradients that may resonate with the color-sensitive filters in pretrained convolutional and transformer encoders.
@@
3.3 Model Configuration
Рис. 6: Different preprocessing methods utilized in experiments. The figure presents viridis colormap, grayscale, and magma colormap.
@@ 3.4 Prompting Strategies
Prompt design is an important since it defines how the model links thermal image features with semantic meaning. Two classes were defined: present, where an object is on the build plate, and absent, where the build plate is empty. Several sentences that describe the same scene in different ways were composed to form the prompt bank for each class. In this bank, prompts for the present class emphasize a clear dark object resting on the heated surface, while prompts for the absent class focus on the surface being empty. For instance, prompts such as “an infrared image showing a bright surface with a solid shape on it” or “a thermal infrared image of a bright surface occupied by a solid object” were used to describe the present class. For the absent class, prompts like “infrared photo of a flat bright surface that is completely empty” or “a thermal image of a bright platform with no blocks or parts” were used. To keep the experiments robust to different phrasings, the same prompt banks were reused across preprocessing methods with only minor wording changes. For magma preprocessing, descriptions were adjusted to refer to an orange surface instead of bright, while for viridis preprocessing the prompts referred to a green–yellow surface.
@@ 5 Conclusions
This work demonstrates a notable advancement in the field of scene description for manufacturing by presenting VLM-IRIS, a zero-shot framework that adapts VLMs for use with thermal infrared imagery. This work developed and validated a framework that converts infrared data into RGB-compatible images and pairs them with descriptive text prompts, and does object presence detection without any retraining or labeled datasets. VLM-IRIS operates fully in a zero-shot manner, which makes it easy to maintain and reliable over time, and allows it to keep working well even if the equipment or environment changes. This creates new opportunities for using such models in manufacturing, where thermal cameras are already common for process monitoring, fault detection, and inspection.
@@
The experiments on a 3D printer testbed demonstrated that CLIP (ViT-B/32) can effectively interpret thermal scenes when supported by suitable preprocessing and prompt design. The experiments confirmed the importance of the preprocessing method and highlighted the impact of prompt engineering on model predictions. Among the tested approaches, the magma colormap with centroid-based prompting achieved the best results, reaching 100% accuracy under room-temperature conditions. Grayscale mapping provides the most physically accurate and interpretable representation but can miss objects when thermal contrast is weak. Magma delivers the most consistent performance across conditions by introducing structured color gradients that resemble RGB data CLIP was trained on. Viridis also performs well overall but tends to generate false positives in high-temperature scenes due to reflective heat patterns. In terms of the temperature of the setting, moderate contrast and structured color mappings help CLIP form more stable embeddings, while extreme heat gradients add visual noise that reduces accuracy.
@@
As the results confirm that VLMs which are originally designed for natural RGB imagery can be effectively adapted to thermal infrared domains, the framework will be extended to more complex inspection tasks, such as defect detection in additive manufacturing. Moreover, future work will explore fine-tuning these models on thermal datasets[ghiasvand2025few, kermani2025systematic]. This fine-tuning will help the model better learn the characteristics of infrared imagery in industrial environments, which will improve both accuracy and robustness of the framework. In addition, incorporating controlled positive noise into the transformer architecture may help reduce model entropy and simplify classification [esfandiari2025multi]. Finally, explainable-AI techniques can provide insight into what visual cues CLIP relies on when embedding the image and text that helps guide further optimization of the framework.

### v2

@@ 1 Introduction
Artificial intelligence (AI) is increasingly transforming manufacturing, where computer vision technologies play central roles in many tasks such as monitoring processes, detecting defects, and ensuring quality [mozaffar2025tunes, yazdipaz2025robust, safavigerdini2025automated, dashti2025functional, lalehdashti2025constructing]. However, most vision-based systems today depend on large labeled datasets that are built for specific tasks and environments. This reliance makes them hard to adapt in dynamic manufacturing settings, since collecting and labeling data is both time-consuming and labor-intensive[ziad2025advancing]. In addition, many industrial environments are low light or enclosed, which reduces the reliability of standard optical cameras. In enclosed additive manufacturing systems, printed parts can be difficult to distinguish from the build surface when using RGB cameras, especially under poor lighting conditions. Thermal infrared (IR) imaging offers a strong alternative because it does not depend on ambient light and can capture useful contrast from temperature differences between machine surfaces, tools, and parts. Figure 1 illustrates this contrast by comparing RGB and infrared images under dark conditions.
@@
Zero-shot learning (ZSL) offers a promising alternative that enables models to generalize to unseen tasks without the need for retraining [xian2017zero, aghdam2025actalign]. A ZSL model can make predictions on previously unseen scenarios by using additional information such as descriptive attributes or semantic text descriptions. This approach is especially useful in manufacturing, where new parts, materials, or process conditions are introduced frequently and collecting labeled data for each scenario is often time-consuming and impractical. Hua et al. [hua2022zero] introduced a framework for tool wear prediction under non-stationary machining environments. They combined deep learning with causal inference to extract invariant features across conditions. A tensor-based method for zero-shot fault diagnosis in smart process manufacturing was also developed to evolve dual prototypes to align sample and attribute representations [ren2024zero]. They improved generalization to unseen faults across multiple datasets. Huang et al. [huang2023simple] presented a simple framework for generalized zero-shot fault diagnosis, using binary classifiers with semantic correction to reduce bias in fault attributes. These studies highlight the strength of ZSL in manufacturing because collecting and labeling large datasets is both costly and often impractical.
@@
As discussed above, VLMs have recently shown promise in manufacturing, but their use in this domain is still at an early stage [zhao2025industrial]. A key limitation for utilizing these models in manufacturing tasks is that they are trained almost entirely on natural RGB images [faysal2024nmformer]. This makes it unclear how well they can transfer to other modalities such as thermal infrared imagery [borazjani2025multi]. This is an important gap because thermal cameras are already common in low-light or enclosed industrial environments where RGB cameras are unreliable. However, most current industrial implementations rely on supervised pipelines that are customized to individual machines and processes, and large available infrared datasets with semantic labels or descriptive text are still rare. As a result, many existing approaches require custom data collection and retraining for each new task, which limits how easily they can be scaled or adapted. This gap motivates the use of zero-shot and label-free approaches that can build on existing pretrained models without the need for extensive infrared training data. Some studies in other domains have adapted vision foundation models to infrared data, but no prior work has systematically investigated zero-shot classification with VLMs on thermal IR data in manufacturing tasks.
@@
Reliable object presence detection on the build platform is an important component of autonomous manufacturing workflows. In additive manufacturing environments with downstream automation, such as robotic pick-and-place or transfer to subsequent processing stages, confirming that a printed part is present on the build plate is a necessary prerequisite for safe and reliable operation. Establishing this capability using thermal infrared sensing and zero-shot vision–language models provides a flexible foundation for automation without requiring task-specific retraining. This work introduces VLM-IRIS (Vision–Language Models for InfraRed Industrial Sensing), a zero-shot framework for adapting VLMs to thermal infrared imagery. It combines the strengths of zero-shot VLMs and infrared imaging to enable reliable object presence detection without any model retraining. The framework converts the infrared images into RGB representations suitable for VLMs to bridge the modality gap. It also utilizes a centroid-based prompting strategy, which multiple descriptive sentences are averaged to create stable text embeddings that reduce sensitivity to prompt wording in order to capture the highest similarity between image-text pairs in the VLMs’ semantic shared space. The goal of this work is to better understand how vision–language models can be used with thermal infrared data in manufacturing settings. Specifically, we study whether pretrained vision–language models can be applied to infrared images without retraining, how different infrared-to-RGB preprocessing choices influence zero-shot performance, and how prompt design strategies affect the robustness of infrared scene classification. In particular, we compare preprocessing approaches that preserve the original thermal structure with those that introduce color variations commonly used in scientific visualization, to examine how these choices interact with vision–language models trained on natural RGB imagery. Through these investigations, this work evaluates both the potential and the current limitations of using vision–language models for label-free infrared monitoring in additive manufacturing. The best-performing configuration of the framework combines CLIP with magma preprocessing and centroid-based prompting to achieve strong accuracy in zero-shot classification. To further validate the robustness of VLM-IRIS, we also evaluate alternative settings across different preprocessing pipelines, prompting strategies, and dataset conditions.
@@ 2.3 Embeddings and Classification via Cosine Similarity
After preprocessing the infrared images and defining the prompt banks, a pretrained VLM was utilized to project both the RGB images and the prompts into a shared multimodal embedding space. The model employs two separate encoders: one for images and one for text. The image encoder transforms each infrared image into a feature vector while the text encoder maps each prompt into the same space. All vectors were normalized to unit length for evaluation based on their directions.
@@ 3 Experimental Setup
Рис. 3: Experimental setup featuring a FLIR Boson thermal camera and a Prusa MK3S 3D printer on the table.
@@ 3.1 Data Collection
The Prusa MK3S 3D printer was used to test the framework as a representative manufacturing platform. A FLIR Boson thermal infrared camera was mounted above the printer’s build plate to capture top-down images of the heated bed after printing and provide a consistent view across all experiments. The positioning of the infrared camera can influence the captured thermal patterns, particularly in enclosed build chambers where reflections and occlusions may occur. Since the task focuses on relative thermal contrast for binary object presence detection, no explicit radiometric or geometric calibration was required. Infrared images were captured after print completion to evaluate object presence under controlled thermal conditions. Figure 3 shows the experimental setup that was used for data collection. All test parts were fabricated using PETG filament with the nozzle temperature set to $\sim$230 °C and the bed temperature maintained between 85–110 °C. The collected dataset consists of two classes: object present and object absent. To evaluate the effect of thermal contrast between the surface and the printed part on predictions, data points were captured under a hot build plate ($\sim$85°C) and a room temperature build plate ($\sim$34°C). In total, 200 images were captured, with a balanced distribution across the two categories and temperature conditions. For each printed part, infrared images were captured both immediately after printing on a hot build plate and again after cooling to room temperature. This was done to ensure that the same object geometries were represented under both temperature conditions. To further evaluate generalization, objects with different geometries, shapes, and sizes were printed. The printed parts were intentionally chosen with different shapes and sizes, ranging roughly from 50 to 150 mm in maximum planar extent and 4 to 70 mm in height, to expose the method to a variety of thermal appearances and evaluate how robust the zero-shot object presence detection is across different configurations. Figure 4 shows examples of the printed objects used in this study.
Рис. 4: A few 3D printed parts during the experiments. This figure presents the variety of the shapes and geometries of the printed parts.
@@ 3.2 Preprocessing Variants
In grayscale mapping, the single thermal intensity channel is normalized and replicated across three channels to preserve the true temperature contrast without adding artificial colors. This method serves as a baseline to test whether VLMs can interpret thermal data that is visually simple yet physically accurate. Magma and viridis colormaps were chosen because they are widely used in scientific visualizations. These colormaps transform intensity values into structured RGB color distributions, where pixel intensity values are mapped to colors. The applied colormaps directly encode the measured temperature values by mapping temperature to color and do not introduce additional information beyond the underlying thermal signal. Such mappings can introduce artificial edges and gradients that may resonate with the color-sensitive filters in pretrained convolutional and transformer encoders.
@@
Рис. 5: Different preprocessing methods utilized in experiments. The figure presents viridis colormap, grayscale, and magma colormap.
3.3 Model Configuration
@@
Рис. 6: Example of image-text pairs in shared embedding space. Infrared images of an empty and an occupied 3D printer are projected alongside their textual prompts, showing how the alignment of vectors reflects semantic similarity.
@@ 3.4 Prompting Strategies
In VLMs, images and text are mapped into a shared embedding space where semantic similarity can be directly measured. Each infrared image is encoded into a feature vector and compared against feature vectors generated from text prompts that describe possible scene interpretations. The prompt whose embedding is most similar to the image embedding determines the predicted class. Figure 6 illustrates this process in a simplified two-dimensional embedding space. VLMs work by matching what is seen in an image with what is described in text. In this context, prompts written by humans are used to clearly describe the task using visual cues that people naturally notice such as contrast, color, and whether a surface is occupied by an object. More descriptive and task-relevant prompts help guide the model toward relevant visual patterns. Prompt design is important since it defines how the model links thermal image features with semantic meaning. Because different word choices can influence the outcome, this work employs a prompt bank, where embeddings from multiple prompts are averaged to form a more stable representation for each class and reduce sensitivity to individual wording. Two classes were defined: present, where an object is on the build plate, and absent, where the build plate is empty. Several sentences that describe the same scene in different ways were composed to form the prompt bank for each class. In this bank, prompts for the present class emphasize a clear dark object resting on the heated surface, while prompts for the absent class focus on the surface being empty. For instance, prompts such as “an infrared image showing a bright surface with a solid shape on it” or “a thermal infrared image of a bright surface occupied by a solid object” were used to describe the present class. For the absent class, prompts like “infrared photo of a flat bright surface that is completely empty” or “a thermal image of a bright platform with no blocks or parts” were used. To keep the experiments robust to different phrasings, the same prompt banks were reused across preprocessing methods with only minor wording changes. For magma preprocessing, descriptions were adjusted to refer to an orange surface instead of bright, while for viridis preprocessing the prompts referred to a green–yellow surface.
@@ 5 Conclusions
This work demonstrates a notable advancement in the field of scene description for manufacturing by presenting VLM-IRIS, a zero-shot framework that adapts VLMs for use with thermal infrared imagery. This work developed and validated a framework that converts infrared data into RGB-compatible images and pairs them with descriptive text prompts, and does object presence detection without any retraining or labeled datasets. VLM-IRIS operates fully in a zero-shot manner, which makes it easy to maintain and reliable over time, and allows it to keep working well even if the equipment or environment changes. The framework is not tied to a specific material extrusion printer or filament material since it relies on relative thermal contrast and semantic alignment. When applied to different printers or materials, factors including emissivity, surface finish, and operating temperature ranges may affect how the infrared images appear and how prompts should be written. However, the core zero-shot pipeline would remain unchanged, and therefore adapting the method mainly involves updating the prompt descriptions rather than retraining the model. This creates new opportunities for using such models in manufacturing, where thermal cameras are already common for process monitoring, fault detection, and inspection.
@@
The experiments on a 3D printer testbed demonstrated that CLIP (ViT-B/32) can effectively interpret thermal scenes when supported by suitable preprocessing and prompt design. The experiments confirmed the importance of the preprocessing method and highlighted the impact of prompt engineering on model predictions. Among the tested approaches, the magma colormap with centroid-based prompting achieved the best results, reaching 100% accuracy under room-temperature conditions. Grayscale mapping provides the most physically accurate and interpretable representation but can miss objects when thermal contrast is weak. Magma delivers the most consistent performance across conditions by introducing structured color gradients that resemble RGB data CLIP was trained on. Viridis also performs well overall but tends to generate false positives in high-temperature scenes due to reflective heat patterns. In terms of the temperature of the setting, moderate contrast and structured color mappings help CLIP form more stable embeddings, while extreme heat gradients add visual noise that reduces accuracy. Future work will explore how these findings transfer to other VLMs, since differences in architecture, training data, and embedding behavior may influence sensitivity to preprocessing choices and prompt design.
The object presence detection scenario considered in this study is not intended to represent the full complexity of industrial perception tasks. Rather, it was selected to isolate the challenge of adapting thermal infrared imagery to VLMs in a zero-shot setting. The selected scenario allows us to examine the effects of preprocessing and prompt design without introducing additional complexity from multi-class recognition or geometric reasoning. At the same time, object presence detection remains a practically useful capability for downstream automation in manufacturing environments. As the results confirm that VLMs which are originally designed for natural RGB imagery can be effectively adapted to thermal infrared domains, the framework will be extended to more complex inspection tasks, such as defect detection in additive manufacturing. In future extensions, different defect levels, such as small, medium, or large, can be described with text prompts, and the model can identify the most likely category by comparing the infrared image with these prompt descriptions, without retraining the model.
@@
The proposed methodology has the potential to extend beyond material extrusion processes to a range of additive manufacturing modalities where thermal sensing is commonly employed. The robustness of the approach arises from its reliance on relative thermal patterns and semantic reasoning through VLMs, rather than process-specific training or handcrafted features. However, different additive manufacturing processes can show different thermal behavior and surface characteristics, which may affect infrared appearance. Adapting the method to these settings may require adjusting the preprocessing steps or refining the prompts, which is an important direction for future work. Center cropping in the experiments assumes a fixed camera view and a known build-plate region, which was used here to focus on infrared modality adaptation and zero-shot inference without adding localization complexity. In more general settings, the same framework could be extended by evaluating multiple regions or integrating localization mechanisms, which is left for future work. Moreover, future work will explore fine-tuning these models on thermal datasets[ghiasvand2025few, kermani2025systematic]. This fine-tuning will help the model better learn the characteristics of infrared imagery in industrial environments, which will improve both accuracy and robustness of the framework. In addition, incorporating controlled positive noise into the transformer architecture may help reduce model entropy and simplify classification [esfandiari2025multi]. Finally, explainable-AI techniques can provide insight into what visual cues CLIP relies on when embedding the image and text that helps guide further optimization of the framework.
Список литературы